Hi, I'm having trouble with this problem:
Graph the linear equation by finding the x and y intercepts.
y=-4x-3
Thank you so much for your help!!

Answer by jim_thompson5910(35256)

To find the x-intercept, plug in y = 0 and solve for x


y = -4x-3


0 = -4x-3


3 = -4x


3/(-4) = x


-3/4 = x


x = -3/4


The x-intercept is the point (-3/4, 0). In decimal form, the x-intercept is the point (-0.75, 0) since -3/4 = -0.75


--------------------------------------------------------------------------------------------------------------


To find the y-intercept, plug in x = 0 and solve for y


y = -4x-3


y = -4(0)-3


y = 0-3


y = -3


So the y-intercept is the point (0,-3)


--------------------------------------------------------------------------------------------------------------


The x-intercept is (-0.75, 0)
The y-intercept is (0,-3)


Plot the two points (-0.75, 0) and (0,-3). Then draw a straight line through them to graph y = -4x-3